OpenStudy (hellokitty17):

FAN AND MEDAL!!!!!!!!! A golf ball is hit from the ground with an initial velocity of 208 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? 13 secs 21 secs 42 secs 6 secs

OpenStudy (anonymous):

i got this hold on

oregonduck (oregonduck):

Assuming that the acceleration of gravity is 32 ft/s^2, we see that it takes: 208/32 = 6.5 seconds, for the ball to reach maximum height (which occurs when the instantaneous velocity of the ball is zero). Recall that the time between initial launch and maximum height and the time between maximum height and the ball hitting the ground is the same. Thus, the ball also takes 6.5 seconds to return to the ground. In total, the golf ball is in the air for 6.5 + 6.5 = 13 seconds and the answer is (A) 13 secs.

OpenStudy (anonymous):

Assuming that the acceleration of gravity is 32 ft/s^2, we see that it takes: 208/32 = 6.5 seconds, for the ball to reach maximum height (which occurs when the instantaneous velocity of the ball is zero). Recall that the time between initial launch and maximum height and the time between maximum height and the ball hitting the ground is the same. Thus, the ball also takes 6.5 seconds to return to the ground. In total, the golf ball is in the air for 6.5 + 6.5 = 13 seconds and the answer is (A) 13 secs.

OpenStudy (anonymous):

XD

oregonduck (oregonduck):

XD

OpenStudy (hellokitty17):

Sweet thanks!!!